### v2

Abstract
Generating landscape paintings expands the possibilities of artistic creativity and imagination. Traditional landscape painting methods involve using ink or colored ink on rice paper, which requires substantial time and effort. These methods are susceptible to errors and inconsistencies and lack precise control over lines and colors. This paper presents LPGen, a high-fidelity, controllable model for landscape painting generation, introducing a novel multi-modal framework that integrates image prompts into the diffusion model. We extract its edges and contours by computing canny edges from the target landscape image. These, along with natural language text prompts and drawing style references, are fed into the latent diffusion model as conditions. We implement a decoupled cross-attention strategy to ensure compatibility between image and text prompts, facilitating multi-modal image generation. A decoder generates the final image. Quantitative and qualitative analyses demonstrate that our method outperforms existing approaches in landscape painting generation and exceeds the current state-of-the-art. The LPGen network effectively controls the composition and color of landscape paintings, generates more accurate images, and supports further research in deep learning-based landscape painting generation.
keywords:
image generation; decoupled cross-attention; latent diffusion model; controllability;
[…]
1 Introduction
Landscape painting Andrews (1999); Clunas (1997); Conisbee (1996); Way et al. (2001) holds a significant place in traditional art, depicting natural scenes to convey the artist’s aesthetic vision and emotional connection to nature. This genre utilizes tools such as ink, water, paper, and brushes, emphasizing artistic conception, brush and ink intensity variations, balanced composition, and effective use of negative space. With advances in computer technology, the generation of landscape paintings through modern algorithms has become a promising area of research.
The creative process of landscape painting Guo (1999); Sze (1992) involves several intricate stages: outlining, chapping, rubbing, moss-dotting, and coloring, as shown in Figure 1 (a). Each stage requires a high level of skill and experience. Outlining requires precise control of lines, chapping captures the diverse textures of nature, rubbing, and dyeing necessitate meticulous control over the ink’s intensity and moisture, moss-dotting calls for a flexible and natural technique, and coloring requires a keen sense of color harmony. The complexity and precision involved in each stage contribute to the overall challenge of creating landscape paintings. Recent technological advancements, particularly in deep learning, have empowered computers to emulate landscape paintings’ artistic styles and brushwork. These technologies can generate artworks that capture the rich cultural essence of China, preserving traditional art forms and opening new paths for creative expression. This technological innovation supports the preservation and evolution of conventional culture Chen (2012); Williams and Marsh (2014). Research into the generation of landscape paintings is primarily divided into traditional and deep learning-based methods. Traditional methods include non-photorealistic rendering, image-based approaches, and computer-generated simulations. Non-photorealistic rendering Way et al. (2006, 2001); Liu (2015) uses computer graphics to simulate various artistic effects, such as ink diffusion, through specialized algorithms. Image-based methods Elgammal et al. (2017); Wang et al. (2018); Gatys et al. (2016) involve collecting brush stroke texture primitives (BSTP) Yu et al. (2003) from hand-drawn samples and employing them to map multiple layers to create mountain imagery. Despite their sophistication, image-based and non-photorealistic rendering methods are inherently complex, and their analog numerical modeling quality often falls short. Currently, many researchers are using generative adversarial networks (GANs) Goodfellow et al. (2014) to generate landscape paintings. However, due to the distinctive artistic style of landscape painting, GAN-based methods often encounter challenges such as unclear outlines, uncontrollable details, and inconsistent style migration.
{adjustwidth}
[…]
Figure 1: The creative process of landscape painting involves several key steps: outlining, chapping, rubbing, moss-dotting, and coloring. Utilizing the stable diffusion method, landscape paintings can be precisely and flexibly generated by specifying a style reference image, a canny outline image, and appropriate cue words.
By integrating computer vision with generative modeling techniques, such as stable diffusion (SD) Rombach et al. (2022), the artistic effects of landscape painting can be effectively simulated and reproduced Li et al. (2023); Sun et al. (2024). Diffusion models (DMs) have been successfully applied across various domains, including clothing swapping, face generation, dancing, and anime creation, demonstrating their versatility in generating high-quality realistic content from diverse inputs. For instance, diffusion models can produce highly realistic images nearly indistinguishable from actual photographs in face generation. Similarly, in anime creation, they can generate detailed and stylistically consistent images from simple prompts Lee et al. (2023); Kim et al. (2023); Choi et al. (2023); Yang et al. (2023). This method not only preserves the stylistic features of traditional landscape paintings but also allows precise control over the content and layout. This technology has broad applications in the creation of art, the preservation of cultural heritage, and digital display, as illustrated in Figure  1(b).
Composition and color control are crucial factors in landscape painting, significantly influencing the overall effect and quality of the artwork. Previous style migration methods have often yielded unsatisfactory results in generating landscape paintings. To address this, we propose the LPGen framework, which draws inspiration from the fundamental steps of traditional landscape painting. This framework employs the canny edge detection algorithm to generate line drawings as input, a structure controller module to guide the image generation based on the canny outlines, and a style controller module to manage the painting style of the generated image. The main contributions of this work are summarized as follows:
  • $\bullet$
    We propose LPGen, a high-fidelity, controllable model for landscape painting generation. This model introduces a novel multi-modal framework by incorporating image prompts into the diffusion model.
  • $\bullet$
    We construct a comprehensive dataset comprising 2,520 high-resolution images, meticulously categorized into four distinct styles: azure green landscape, golden splendor landscape, ink wash landscape, and light vermilion landscape.
  • $\bullet$
    We conduct extensive qualitative and quantitative analyses of our proposed model, LPGen, providing clear evidence of its superior performance compared to several state-of-the-art methods.
2 Related Work
2.1 GAN-Based Method
Due to the rapid advancements in deep learning, significant progress has been made in various visual tasks Shen et al. (2023a, b). Based on the input content, landscape painting generation methods can be categorized into three types: nothing-to-image, image-to-image, and text-to-image.
For the first type, it does not require any input information to generate landscape paintings. An example is the Sketch-And-Paint GAN (SAPGAN) Xue (2021), the first neural network model capable of automatically generating traditional landscape paintings from start to finish. SAPGAN consists of two GANs: SketchGAN for generating edge maps and PaintGAN for converting these edge maps into paintings. Another example is an automated creation system Luo et al. (2022) based on GANs for landscape paintings, which comprises three cascading modules: generation, scaling, and super-resolution.
For the second type, a photo is input for style transfer. An interactive generation method Zhou et al. (2019) allows users to create landscape paintings by simply outlining with essential lines, which are then processed through a recurring adversarial generative network (RGAN) model to generate the final image. Another approach, neural abstraction style transfer Li et al. (2019), leverages the MXDoG filter and three fully differentiable loss terms. The ChipGAN architecture He et al. (2018) is an end-to-end GAN model that addresses critical techniques in ink painting, such as blanks, brush strokes, ink tone and spread. Additionally, an image style transfer framework Peng et al. (2022) utilizing CycleGAN, facilitates the conversion of photographs into paintings. The bilateral generative adversarial network algorithm (BiTGAN) He et al. (2023) constructs a bilateral generator using a shared encoder, U-Net path, ResNet path and decoder, incorporating an attentional network Shen et al. (2023) to convert photos into paintings. Furthermore, an attentional wavelet network Wang et al. (2021) utilizes wavelets to capture high-level mood and local details of paintings via the 2-D Haar wavelet transform. Lastly, the Paint-CUT model Wang et al. (2023) intelligently creates landscape paintings by utilizing a shuffle attentional residual block and edge enhancement techniques.
For the third type, the output image is guided by textual input information. A novel system Yuan et al. (2023) for Polaca transforms classical poetry into corresponding artistic landscape paintings and calligraphic works. Another approach, controlled landscape painting generation (CCLAP) Wang et al. (2023), is based on the latent diffusion model (LDM) Ho et al. (2020) and comprises two sequential modules: a content generator and a style aggregator. Although this method allows text-based control of the image, it still faces challenges related to the randomness of image generation and limited controllability.
2.2 Diffusion-Based Method
DM Ho et al. (2020) is a deep learning framework primarily utilized for image processing and computer vision tasks. Convert an input image into Gaussian noise through diffusion steps, and then reconstruct the original image via an inverse process. DM can generate detailed images from textual descriptions and are also effective for image restoration, image drawing, text-to-image, and image-to-image tasks Chen et al. (2024); Lian et al. (2024); Shen et al. (2023, 2024). Essentially, by providing a textual description of the desired image, SD can produce a realistic image that aligns with the given description Ho et al. (2020). These models reframe the "image generation" process into a "diffusion" process that incrementally removes noise. Starting with random Gaussian noise, the models progressively eliminate it through training until it is entirely removed, yielding an image that closely matches the text description Ho et al. (2020); Lian et al. (2024). However, a significant drawback of this approach is the considerable time and memory required, especially for high-resolution image generation Chen et al. (2024). LDM Rombach et al. (2022) was developed to address these limitations by significantly reducing memory and computational costs. This is achieved by applying the diffusion process within a lower-dimensional latent space rather than the high-dimensional pixel space Chen et al. (2024).
The emergence of large text-to-image models that can create visually appealing images from brief descriptive prompts has highlighted the remarkable potential of artificial intelligence. However, these models encounter challenges such as uneven data availability across specific domains compared to the generalized image-to-text domain. In addition, some tasks require more precise control and guidance than simple prompts can provide. ControlNet Zhang et al. (2023) addresses these issues by generating high-quality images based on user-provided cues and controls, which can fine-tune performance for specific tasks. Similarly, Ip-adapter Ye et al. (2023) employs decoupled cross-attention mechanisms for the characteristics of the text and the image. Inspired by these approaches, our study incorporates ControlNet’s additional structural controls, such as sketches, depth maps, and semantic segmentation maps. Moreover, the Ip-adapter can utilize style or content from reference images to guide the generation of landscape paintings. For example, the Thangka sketch colorization method Li et al. (2023) proposed by Li et al. employs multi-level adaptive instance-normalized color fusion combined with skip connection attention to effectively capture and distinguish the intricate semantic information embedded within the sketch lines. Additionally, Edge-MSGAN Sun et al. (2024) harnesses edge images to guide a completion network in generating complete ancient paintings. It employs a multibranch color correction network to precisely adjust and harmonize the colors, ensuring fidelity to the original artistic style.
3 Proposed Method
To address the above issues, LPGen has been developed to generate landscape paintings from photos. This section provides a comprehensive introduction to LPGen, detailing an overview of the model, the image structure controller, which manages the composition of the generated image, and the style controller, which dictates the painting style of the resulting image.
{adjustwidth}
[…]
Figure 2: Our proposed LPGen framework for landscape painting generation comprises three key components. First, the image structure controller module utilizes the Decoupled Cross-Attention technique to separately manage the structural and stylistic information of an image across different domains, allowing for precise control and regulation of specific elements and attributes in the generated image. Second, the style controller dynamically adjusts the features of the input image, enabling the generative model to accurately capture and reflect the style and structure of the source image. Finally, the text expression controller guides the image generation model using natural language descriptions, facilitating the production of high-quality images that align with the provided textual prompts.
3.1 Preliminaries
The complete structure of the LPGen framework is illustrated in Figure 2. The framework comprises three main components: the stable diffusion model, the structure controller, and the style controller. While the stable diffusion model offers generalized image generation capabilities, it lacks precise control over the structural and stylistic features of an image. To address this, we employ the structure controller to ensure that edge information effectively guides the image generation process. This controller guarantees that the generated image aligns with the composition of the canny reference image. Artists can dictate the position and shape of graphic elements through outlining, and LPGen uses the structure controller to manage the structure and form of the resulting image. For stylistic aspects, the style controller extracts the style of a reference segmented image and applies it to the generated image, allowing the generation of images in a specified style. The style controller learns various color schemes to convey different emotions. Building upon the stable diffusion model’s base image generation capabilities, the structure controller manages the structural layout, and the style controller governs the color scheme features. This precise and robust control enables us to achieve excellent management over the style and quality of the generated images, surpassing the capabilities of simple text-to-image models.
3.2 Structure Controller
The line drawing serves as the fundamental framework of landscape painting, establishing the overall layout and positioning of elements. During the coloring stage, the artist sequentially fills and renders the picture, coloring each element according to the guidelines set by the line drawing. To enable the line drawing to guide LDM in generating the image, the structure controller manipulates the neural network structure of the diffusion model by incorporating additional conditions. This, in combination with Stable Diffusion, ensures accurate spatial control, effectively addressing the spatial consistency issue. The image generation process thus emulates the actual painting process, transitioning from line drawing to coloring. We employ the structure controller as a model capable of adding spatial control to a pre-trained diffusion model beyond basic textual prompts. This controller integrates the UNet architecture from Stable Diffusion with a trainable UNet replica. This replica includes zero convolution layers within the encoder and middle blocks. The complete process executed by the structure controller is as follows:
$y_{c}=\mathcal{F}(x,\theta)+\mathcal{Z}(\mathcal{F}(x++\mathcal{Z}(c,\theta_{z1}),\theta_{c}),\theta_{z2}).$ (1)
The structure controller differentiates itself from the original SD in handling the residual component. $\mathcal{F}$ signifies the UNet architecture, with $x$ as the latent variable. The fixed weights of the pre-trained model are denoted by $\theta$. Zero convolutions are represented by $\mathcal{Z}$, having weights $\theta_{z1}$ and $\theta_{z2}$, while $\theta_{c}$ indicates the trainable weights unique to the structure controller. Essentially, the structure controller encodes spatial condition information, such as that from canny edge detection, by incorporating residuals into the UNet block and subsequently embedding this information into the original network.
3.3 Style Controller
In the original SD base model with text-generated graph capability, we introduce the style controller structure to integrate text cueing features, structure cueing features and image style cueing features. To achieve the above purpose, our core design is a decoupled cross-attention mechanism as shown in Figure 2 , where image features are embedded through a newly added cross-attention layer, which consists mainly of two modules: an image encoder for extracting image features, and an adaptation module with decoupled cross-attention for embedding the image features into a pre-trained text-to-image diffusion model.
The pre-trained CLIP image encoder model is used, but here, in order to efficiently decompose the global image Embedding, we use a small trainable projection network to project the image Embedding into a sequence of features of length N=4, which is the Image Feature in Fig. The projection network here is designed as a linear layer Linear plus a normalization layer LN, and at the same time, the dimensions of the input image features are kept consistent with the dimensions of the text features in the pre-trained diffusion model.
In the original Stable Diffusion model, text embeddings are injected into the Unet model through the input-to-cross-attention mechanism. A straightforward way to inject image features into the Unet model is to join image features with text features and then feed them together into the cross-attention layer. However, this approach is not effective enough. Instead, the style controller separates the cross-attention layer for text features and image features by decoupling this mechanism, making the model more concise and efficient. This design not only reduces the demand for computational resources, but also improves the generality and scalability of the model. During the training process, the style controller is able to automatically learn how to generate corresponding images based on text descriptions, while maintaining the effective utilization of image features. This enables style controller to generate images with full consideration of the semantic information of the text, thus generating more accurate and realistic images.
The text corresponds to the cross-attention as:
$\mathcal{Z}_{new}=\mathcal{\textit{Attention}}(Q,K^{t},V^{t})+\lambda\cdot\mathcal{\textit{Attention}}(Q,K^{i},V^{i}),$ (2)
Here, $Q$, $K^{t}$, and $V^{t}$ represent the query, key, and value matrices for the text cross-attention operation, while $K^{i}$ and $V^{i}$ are the key and value matrices for the image cross-attention. Given the query features $Z$ and the image features $c_{i}$, the formulations are as follows: $Q=ZW_{q}$, $K^{i}=c_{i}W_{k}^{i}$, and $V^{i}=c_{i}W_{v}^{i}$. It is important to note that only $W_{k}^{i}$ and $W_{v}^{i}$ are trainable weights.
3.4 Training and Inference
During training, we focus solely on optimizing the style controller, leaving the parameters of the pre-trained diffusion model unaltered. The style controller is trained using a dataset of paired images and text. Still, it can train without text prompts, as only image prompts effectively guides the final generation. The training objective remains consistent with that of the original SD:
$L_{\text{simple}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t}\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big{(}\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t\big{)}\|^{2}.$ (3)
During the training phase, we consistently employ random omission of image conditions to facilitate classifier-free guidance during inference.
$\begin{split}\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)=w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)+(1-w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t),\end{split}$ (4)
If the image condition is omitted, we replace the CLIP image embedding with a zero vector.
Since text cross-attention and image cross-attention are separate, we can independently adjust the weight of the image condition during inference.
$\mathbf{Z}^{new}=\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})+\lambda\cdot\text{Attention}(\mathbf{Q},\mathbf{K}^{\prime},\mathbf{V}^{\prime}),$ (5)
Here, $\lambda$ is a weighting factor. When $\lambda=0$, the model defaults to the original text-to-image diffusion model.
4 Experiment and Analysis
4.1 Datasets
{adjustwidth}
[…]
Figure 3: In this study, data processing involved the following four steps: (1) Collecting raw data, including images and text sources. (2) Cleaning the data to eliminate noise, duplicates, and irrelevant information. (3) Preprocessing of data, such as resizing images and converting formats. (4) Creating matching pairs of text and images for model training.
Limited datasets are available for landscape paintings. The traditional landscape painting dataset collected by Xue  Xue (2021) holds significant value. Building upon this dataset, we have expanded our collection to include various styles of landscape paintings and standardized the corresponding text descriptions. This effort aims to contribute positively to the research on landscape painting generation, as illustrated in Figure 3.
Collection: Our investigation revealed a scarcity of landscape painting datasets available for research. Currently, the Traditional Landscape Painting Dataset Xue (2021) is the only accessible dataset, but it has limitations in terms of data volume and lacks categorization or identification of the image data. To supplement this, we have curated a collection of 500 digital paintings sourced through websites such as Google and Baidu, artwork websites, photographs from landscape painting books, and digital museum databases. The size distribution of these samples is detailed in Table 1 below. To ensure representativeness, we included works from various periods and historical stages.
Table 1: This table is based on the minimum images table for statistical information. Images with small edges greater than or equal to 512 pixels are our effective data and can be used for model training in the future.
Category Count
Less than 512 319
Equal to 512 2617
Greater than or equal to 512 16
Cleaning: We enlisted the help of professors and landscape painters to assist in the image-cleaning process. To ensure the dataset’s quality, these experts manually removed paintings that did not depict landscapes and those with dimensions smaller than 512 pixels or with unclear imagery. Landscape painting is a comprehensive art form that integrates poetry, calligraphy, seal cutting, and painting, unique aspects of this artistic tradition. Directly using these paintings as training data can result in the model generating images with inexplicable elements or words. To better emphasize the theme and beauty of landscape paintings while removing the interference of calligraphy and seal engraving, we utilized image processing software like Adobe Photoshop to repair the damaged area, enhancing image clarity and quality. We adjusted the brightness, contrast, hue, and saturation parameters to make the images sharper and brighter and to highlight their details and features. In addition, we manually used the repair tool to meticulously restore the integrity and beauty of the original images, employing multiple overlapping repair pen strokes to achieve a more complete restoration.
Preprocessing: Since the training data for this study should be in $512\times 512$ format, I processed the images accordingly. First, each image was scaled so that its shorter side was 512 pixels, maintaining the original aspect ratio. For images with an aspect ratio less than 1.5, a $512\times 512$ section was cropped from the center and saved. For images with an aspect ratio greater than 1.5, after cropping the $512\times 512$ section from the center, the center point was moved 256 pixels in both directions along the longer side to crop and save two additional images. If the cropped image was not square, the process was halted, and the current cropped image was discarded.
Text-image Pair Generation: We enlisted landscape painters and professors to categorize the previously processed photos and the Traditional Landscape Painting Dataset. The categories identified are azure green landscape, golden splendor landscape, ink wash landscape, and light vermilion landscape. To facilitate successful model training, the data format for this experiment is image-text pairs. Although image-text pairs previously required manual annotation, we now utilize BLIP-2 to automatically generate the corresponding text information for the collected images. Although the text generated by BLIP-2  Li et al. (2022) may be inaccurate, we invited professional teachers to refine and augment the text descriptions specific to landscape paintings. To increase the diversity of text descriptions, we provided multiple expressions for each text and specified the painting type within the descriptions. Finally, we saved the image paths and corresponding text descriptions in JSON files. This newly curated dataset, comprising the base data and the additional image and image-text pairs, was used to train the new model.
4.2 Evaluation Metrics
Several metrics are commonly used to assess how closely generated images resemble reference images in evaluating image and structural similarity. This document explains the meanings, formulas and applications of six key metrics: the learned perceptual image patch similarity (LPIPS) Zhang et al. (2018), gram matrix Gatys et al. (2015), histogram similarity Swain and Ballard (1991), chamfer match score Barrow and Tenenbaum (1977), hausdorff distance Huttenlocher et al. (1993), and contour match score Belongie et al. (2002).
Learned Perceptual Image Patch Similarity: LPIPS is a metric to evaluate image similarity. It utilizes feature representations of deep learning models to measure the perceptual similarity between two images. Unlike traditional pixel-based similarity metrics (e.g., mean square error or peak signal-to-noise ratio), LPIPS focuses more on the perceptual quality of the image, i.e., the perception of image similarity by the human visual system. The formula for LPIPS is given by:
$\text{LPIPS}(I_{1},I_{2})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\|w_{l}\cdot(\phi_{l}(I_{1})_{h,w}-\phi_{l}(I_{2})_{h,w})\|_{2}^{2},$ (6)
where: $\phi_{l}$ denotes the feature representation at layer $l$. $H_{l}$ and $W_{l}$ are the height and width of the feature map on layer $l$. $w_{l}$ is a learned weight at layer $l$. $I_{1}$ and $I_{2}$ are the two compared images.
Gram Matrix: The gram matrix is a metric for evaluating the stylistic similarity of images and is commonly used in style migration tasks. It captures the texture information and stylistic features of an image by computing the inner product between the feature maps of a convolutional neural network. Specifically, the gram matrix describes the correlation between different channels in the feature map, thus reflecting the overall texture structure of the image.
The formula for the gram matrix at a particular layer is:
$G_{ij}^{l}=\sum_{k}F_{ik}^{l}F_{jk}^{l},$ (7)
where: $F_{ik}^{l}$ is the activation of the $i$-th channel at layer $l$. The sum $\sum_{k}$ is taken over all spatial locations of the feature map.
Histogram Similarity (Bhattacharyya Distance): The histogram similarity is a metric for evaluating the similarity of color distributions of two images and is widely used in image processing and computer vision. By comparing the color histograms of the images, the similarity of the images in terms of color can be determined.The bhattacharyya Distance  Bhattacharyya (1943) is a commonly used histogram similarity metric, especially for similarity calculation of probability distributions.
The formula for the bhattacharyya distance is:
$D_{B}(p,q)=-\ln\left(\sum_{x\in X}\sqrt{p(x)q(x)}\right),$ (8)
where: Let $p(x)$ and $q(x)$ represent the probability distributions of the two histograms to be compared. The summation $\sum_{x\in X}$ is taken in all the bins of the histograms.
Chamfer Match Score: The chamfer match score is a metric for evaluating the similarity of two sets of point clouds, widely used in computer vision and graphics. It quantifies the shape similarity between two sets of point clouds by calculating the average nearest distance between them. Specifically, chamfer match score is a variant of Chamfer Distance and is commonly used for tasks such as 3D shape matching, image alignment and contour alignment.
The formula for the chamfer distance is:
$d_{\text{Chamfer}}(A,B)=\frac{1}{|A|}\sum_{a\in A}\min_{b\in B}\|a-b\|+\frac{1}{|B|}\sum_{b\in B}\min_{a\in A}\|a-b\|,$ (9)
where: $A$ and $B$ represent the sets of edge points in two images. $\|a-b\|$ denotes the Euclidean distance between points $a$ and $b$.
{adjustwidth}
[…]
Figure 4: This figure showcases an example of landscape painting generated using our method. Our approach is highly responsive to the canny edge maps and style reference images. Columns (a), (b), (c), and (d) represent landscape paintings generated in the same style but with different canny edge maps. Conversely, rows (e), (f), (g), and (h) depict landscape paintings generated with the same canny edge map but using different style references.
Hausdorff Distance: The hausdorff distance is a metric for evaluating the maximum distance between two sets of point sets (point clouds, contours, etc.) and is widely used in computer vision, graphics, and pattern recognition. It measures the distance between the farthest corresponding points in two point sets, reflecting the degree to which they differ geometrically.
The formula for the hausdorff distance is:
$d_{H}(A,B)=\max\left\{\sup_{a\in A}\inf_{b\in B}\|a-b\|,\sup_{b\in B}\inf_{a\in A}\|a-b\|\right\},$ (10)
where: $A$ and $B$ denote the sets of edge points in the two images. $\|a-b\|$ represents the Euclidean distance between the points $a$ and $b$. $\sup$ signifies the supremum (the least upper bound) and $\inf$ indicates the infimum (the highest lower bound).
Contour Match Score: The contour match score evaluates the similarity between the contour shapes of two images, making it particularly useful in applications where the overall shape and structure of objects are crucial. This score is determined by comparing the shape descriptors of the contours in the images.
The formula for the contour match score is:
$d_{\text{Contour}}(A,B)=\sum_{i=1}^{n}\left(\frac{(A_{i}-B_{i})^{2}}{A_{i}+B_{i}}\right),$ (11)
where: $A_{i}$ and $B_{i}$ represent the shape descriptors of the contours in the two images. The summation is taken over all contour points $i$.
4.3 Experimental Results
As shown in Figure 4, we use the LPen model to generate landscape paintings. This model learns the characteristics and styles of numerous classic landscape paintings and automatically creates new, highly realistic works. The generated landscape paintings feature typical natural elements such as mountains, rivers, trees, and stones and retain the ink color and brush techniques characteristic of traditional landscape paintings. Additionally, the LPen model can generate new works through innovation and variation.
The LPen model effectively and accurately controls image generation through edge detection and stylized reference images. Using the canny edge detection method, the LPen model extracts edge information from the input image, preserving and improving structural features during the generation process. Meanwhile, the style reference image provides detailed information about the target style, enabling the LPen model to learn and apply these features, thereby aligning the generated image with the expected visual effect. By combining these two control methods, the LPen model generates highly realistic landscape paintings and flexibly adjusts the generated images’ style and details, achieving precise control and innovative performance in traditional art creation.
4.4 Qualitative Analysis
To rigorously assess the effectiveness of the proposed LPGen model in generating high-quality landscape paintings, we conduct a comprehensive comparative analysis against several benchmark methods, including Reference Only Zhang et al. (2023), Double ControlNet Zhang et al. (2023), and Lora Hu et al. (2021).
The comparison experiment with other methods for generating landscape paintings is shown in Figure 5. In this figure, Figure 6a displays the constraint canny edge map, while Figure 6b shows the target ink-style reference image. Figure 6c presents the ink-style image generated by the Lora method. Here, the generated image exhibits noticeable noise and incorrect image information and is limited to Lora’s inherent style, failing to replicate the reference image’s style. Figure 6d illustrates the image produced by the Reference Only method. This image has a more precise contour structure but introduces redundant contours, lacks distinct style characteristics, and suffers from color inaccuracies. Figure 6e shows the ink-style image generated by the Double ControlNet method. Although the constraint contours are well-defined, the style characteristics of the reference image are not effectively learned.Finally, Figure 6f displays the images generated by LPGen.
In summary, the proposed LPGen effectively addresses significant issues such as detail loss, poor style transfer, and blurred lines in generated images, resulting in high-quality landscape paintings. As illustrated in Figure 5, our model consistently outperforms other state-of-the-art methods. LPGen not only preserves the layout and content of the target photos but also successfully captures the essence of the ink wash style, thereby achieving superior overall quality and fidelity.
{adjustwidth}
[…]
Figure 5: Samples of landscape paintings generated using different methods are shown below. Column (a) displays the canny edge image. Column (b) shows the reference image. Column (c) presents the image generated by Lora. Column (d) features the image generated by Reference Only. Column (e) illustrates the image generated by Double ControlNet. Finally, column (f) displays the image generated by our LPGen model.
4.5 Quantitative analysis
The results depicted in Figure 5 align with the quantitative metrics presented in Table 2. The LPIPS, gram matrix, and histogram similarity metrics provide robust and complementary measures to evaluate image similarity. LPIPS assesses perceptual similarity by leveraging deep feature representations, while the gram matrix evaluates style similarity by capturing texture correlations. Histogram similarity, on the other hand, compares color distributions by analyzing histograms. Together, these metrics form a comprehensive framework for determining how closely generated images resemble reference images in both structure and style. Additionally, the chamfer match score, hausdorff distance, and contour match score offer robust and complementary measures for evaluating edge and contour similarity. The chamfer match score focuses on the average similarity of edges, the hausdorff distance emphasizes the worst-case edge similarity, and the contour match score evaluates overall shape similarity. Collectively, these metrics provide a detailed framework for assessing how closely generated images match reference images in terms of edges and contours.
Table 2: The table compares style and structure metrics for images generated by different models. LPIPS evaluates perceptual similarity, GM (gram matrix) measures texture correlations for style similarity, and HS (histogram similarity) assesses color distribution. CMS i(chamfer match score) focuses on average edge similarity, HD (hausdorff distance) emphasizes worst-case edge similarity, and CMS ii(contour match score) evaluates overall shape similarity.
{adjustwidth}
[…]
The first three metrics, LPIPS, the gram matrix, and histogram similarity, are utilized to analyze the structural similarity between the model outputs and the reference images. LPIPS (Learned Perceptual Image Patch Similarity) measures perceptual similarity between images, where lower LPIPS values indicate higher similarity in terms of human visual perception. According to the table, Lora (0.52) performs best in terms of LPIPS, indicating that Lora’s generated images are most perceptually similar to the reference image. The gram matrix metric measures style similarity, particularly in texture, with lower values indicating higher style similarity. The table shows that LPGen (3.40e-06) performs best regarding the gram matrix, suggesting that LPGen’s generated images have the most similar texture style to the reference image. Histogram similarity, measured using the bhattacharyya distance, evaluates the similarity of color distributions, with lower values indicating more identical color distributions. LPGen (0.72) performs best regarding histogram similarity, suggesting that LPGen-generated images have the most similar color distribution to the reference image. For structural similarity, LPGen excels in both texture style (gram matrix) and color distribution (histogram similarity), while Lora performs best in perceptual similarity (LPIPS) due to the structure controller. However, Lora requires repetitive fine-tuning and is computationally intensive. Taking these factors into account, LPGen is the best model for overall structural similarity.
The last three metrics, the chamfer match score, hausdorff distance, and contour match score, are used to analyze the style similarity between different model outputs and the reference image, mainly focusing on edges and contours. The chamfer match score measures edge similarity, where higher scores indicate more remarkable similarity. LPGen (-0.15) performs best with respect to the chamfer match score, suggesting that LPGen’s generated images have the most similar edges to the reference image. The hausdorff distance measures the maximum edge distance between two images, with lower values indicating more remarkable similarity. LPGen (154.54) performs best regarding hausdorff distance, suggesting that LPGen’s generated images are closest to the reference image regarding edge similarity. The contour match score measures the similarity of contour shapes, with lower values indicating more remarkable similarity. LPGen (3.24) performs best regarding the contour match score, indicating that LPGen’s generated images have the most similar contour shapes to the reference image. LPGen outperforms other models in style similarity in all three metrics: chamfer match score, hausdorff distance, and contour match score. This demonstrates that LPGen’s generated images have the most similar edges and contour shapes to the reference image, making it the best model for style similarity.
4.6 Visual Turing Tests
In the user study, we evaluated four critical aspects of the generated images: aesthetic appeal, style consistency, creativity, and detail quality. Aesthetic Appeal: This metric evaluates the overall visual attractiveness of the images. The participants rated the images based on how pleasing they found them, considering factors such as color harmony, composition, and the emotional response evoked by the artwork. Style Consistency: This aspect examines how well the generated images adhere to a specific artistic style. Participants evaluated whether the images consistently incorporated stylistic elements of traditional landscape paintings, such as brushstroke techniques, use of space, and traditional motifs. Creativity: This aspect measures the originality and innovation of the generated images. The participants rated the images based on the novelty and inventiveness of the compositions and interpretations within the confines of traditional landscape painting. Detail Quality: This metric focuses on the precision and clarity of the finer details within the images. Participants evaluated the quality of intricate elements such as textures, line work, and depiction of natural features such as trees, rocks, and water.
Figure 6: The above are sample questions for a user survey. Each question presents four options, one of which is an image generated by this study. The questions evaluate the images’ overall aesthetic appeal, style consistency, creativity, and detail quality.
In our study, we used the same canny and style reference images to generate output with different models. We selected the best options based on four evaluation metrics: aesthetic appeal, style consistency, creativity, and detail quality, as shown in Figure 6. The results in Table 3 from these evaluations provide insights into the effectiveness of the LPen model in generating high-quality, artistically compelling landscape paintings.
Table 3: This table presents the statistical data from the user survey questions. Each data point represents the top-rated results, as determined by users, for images generated by different models in terms of aesthetic appeal, style consistency, creativity, and detail quality.
{adjustwidth}
[…]
In terms of style consistency, LPGen excelled with an impressive score of 61. 46%, far exceeding the Reference Only model (11.46%), the Double ControlNet model (16.66%) and the Lora model (10.42%). This high score indicates that LPGen-generated images consistently adhered to the specific artistic style more effectively, ensuring uniformity in stylistic elements crucial for cohesive and authentic artwork. When it comes to evaluating creativity, LPGen once again led the pack with a top score of 40.63%. This metric highlights the model’s ability to produce novel and inventive compositions within the realm of traditional landscape painting. In contrast, the Reference Only model scored 22.92%, the Double ControlNet model 23.95%, and the Lora model 12.50%. LPGen’s higher score in creativity underscores its capacity to generate images that are not only original but also creatively engaging. Regarding detail quality, LPGen achieved the highest score of 52.08%, showcasing its superiority in rendering intricate elements with precision and clarity. This capability is particularly important for realistically depicting natural features such as trees, rocks, and water. The Reference Only model scored 15.63%, the Double ControlNet model 19.79%, and the Lora model 12.50%. LPGen’s outstanding performance in detail quality highlights its proficiency in producing images with high fidelity and meticulous attention to fine details. Considering these results from all four metrics, the LPGen model’s outstanding performance across these metrics—aesthetic appeal, style consistency, creativity, and detail quality—highlights its effectiveness in generating high-quality, artistically compelling landscape paintings. These results reflect the model’s ability to meet and exceed user expectations in various aspects of image generation.
5 Conclusions
In the conclusion of this paper, we provide a comprehensive examination of our use of quantitative, qualitative, and user-study methodologies to evaluate the performance of various models in generating artistic-style images. Our research combined these approaches to offer a holistic assessment. Quantitative analysis used metrics such as aesthetic appeal, style consistency, creativity, and detail quality to measure model performance. For example, LPGen consistently outperformed other models in aesthetic appeal and detail quality, achieving the highest scores in these categories. Qualitative analysis delved deeper into user evaluations and perceptions of the generated images. This analysis revealed the strengths and weaknesses of different models in replicating artistic styles and expressing creativity. Through qualitative feedback, we gained insights into user experiences and preferences that quantitative metrics alone could not capture. User studies were pivotal in validating our quantitative and qualitative findings. These studies confirmed LPGen’s superiority in visual appeal and innovation from the user perspective, reinforcing the quantitative data and adding context to the qualitative insights.
In summary, our study employed a diverse set of research methodologies to thoroughly evaluate LPGen’s performance in generating artistic-style images. These analyses provided robust quantitative data and deep qualitative insights, enhancing our understanding and advancing the field of artistic generation models. We also noted that LPGen demands substantial computational resources, especially when processing large-scale data or generating high-resolution images. Future research will focus on enhancing the structure controller to more accurately manage generated structures, thereby further improving the quality and diversity of the generated images.
[…]
Conceptualization, W.Y. and Y.Z.; methodology, W.Y.; validation, W.Y., X.W. and Y.Q.; formal analysis, X.W. and Y.Q.; investigation, X.W. and Y.Q.; data curation, X.W. and Y.Q.; writing—original draft preparation, W.Y.; writing—review and editing, Y.Z.; supervision, Y.Z.; project administration, Y.Z.; funding acquisition, W.Y. All authors have read and agreed to the published version of the manuscript.
[…]
This work is jointly supported by the Beijing Municipal High-Level Faculty Development Support Program (Grant No. BPHR202203072).
[…]
The dataset used in this study will be made publicly available at a later stage and can be requested by contacting aiexplorercn@163.com.
Acknowledgements.
[…]